# Irreducibility and periodicity in $\mathbb{Z}^{2}$ symbolic systems

### v1

@@ 2 Coins and buckets (a combinatorial lemma)
Note that when coins are moved from one bucket to another we often do not control the orientation they will receive. In particular, if a coin is added to a non-empty bucket, that bucket cannot be oriented after the move. Observe also that that legal moves do not change the total number of coins in a configuration, and that in each bucket the coins of a given orientation are interchangeable. in the sense that if one of them is to be moved, it does not matter which it is.
@@
Let $c_{k,n}$ denote the configuration consisting of $k$ buckets, all of which are empty except for the first, which contains $n_{1}=n$ coins, out of which $n_{2}=\left\lfloor n_{1}/2\right\rfloor$ are heads and $\left\lceil n_{1}/2\right\rceil$ are tails. Consider the following sequence of moves starting from $c_{k,n}$. First, move $n_{2}$ heads from the first bucket to the second bucket, one by one, orienting them as “tails” whenever there is a choice. Thus, when the first coin lands in an empty bucket its orientation is tails, the second coin is forced to be heads, the next is again tails, and so on. When all $n_{2}$ coins have been transferred, we will have $n_{3}=\left\lfloor n_{2}/2\right\rfloor$ heads and $\left\lceil n_{2}/2\right\rceil$ tails in the second bucket. Now move all heads from the second bucket to the third in the same manner, and continue, ending in the $k$-th bucket. At the end. $\left\lceil n_{i}/2\right\rceil$ tails remain in the buckets $i=1,\ldots,k-1$, and $n_{k}$ coins remain in the last bucket. The final configuration is be oriented if and only if $n_{k}\leq 1$, and, tallying how many coins are left in each bucket have shown that $c_{k,n}$ is orientable if $n\leq 2^{k+1}-1$. Conversely,
@@
We begin the proof. Assume that $k,n$ are given and that $m_{1}\ldots m_{N}$ is a sequence of legal moves taking $c=c_{k,n}$ to an oriented configuration $c^{\prime}$. Our goal is to deduce that $n\leq 2^{k}-1$. Our strategy will be to reduce the given moves to a different sequence, in which each bucket is dealt with separately; then we will be able to directly analyze the situation siilarly to the exmaple above.
@@
To simplify analysis, we first modify the rules of the game. Introduce a special bucket called “the pile”, which serves as a temporary holding place for coins. While in the pile, coins do not have a specific orientation. We only allow moves that transfer a coin from a bucket to the pile, or from the piile to a bucket. The game begins with all coins in the piule, and must end in an empy pile and oriented buckets.
@@
    An original move that transfers a coin from bucket $B$ to bucket $B$’ is replaced by two moves:L one taking a coin from $B$ to the pile, and the other taking a coin from the pile to $B^{\prime}$.
Evidently, starting from $n$ coins on the pile and $k$ empty buckets, the first $n$ moves in the updated sequence resut in the configuration $c$, and the ramining move bring us to $c^{\prime}$.
@@
$\ell_{B}=\#\text{ of coins in }B\text{ in the final configuration }c^{\prime}$
and let $\sigma_{B}$ denote their common orientation (if $B$ is empty in $c^{\prime}$ set $\sigma_{B}=$tails). Let $\sigma^{\prime}_{B}\neq\sigma_{B}$ denote the other orientation, and set
@@
Define the seqeunce of moves $\underline{m}_{B}=m_{B,1},\ldots,m_{B,\ell_{B}+2\ell^{\prime}_{B}}$ as follows:
  • The first $\ell_{B}+\ell^{\prime}_{B}$ moves insert coins into $B$, with th provision that if relevant, the orientation is $\sigma_{B}$.
@@
Then, if we apply $\underline{m}_{B}$ starting from a configuration in which $B$ is empty and there are at least $\ell_{B}+\ell^{\prime}_{B}$ coins in the pile, then the net result is that $\ell_{B}$ coins are transfered to $B$ with orientation $\sigma_{B}$. Note that no shorter sequence of moves can achieve this result.
@@
Let $t_{B}$ denote the maximal integer $1\leq t_{B}\leq N$ such that, after performing the move $m_{t_{B}}$, the bucket is in its peak state, but was not before. Observe that the move at time $t_{B}$, if it exists, must move a coin to $B$, so the times $t_{B}$ are distinct for different buckets, provided that $p_{B}>0$, while if $p_{B}=0$, then the bucket remains throughout, and we can delete all such buckets and reduce $k$. Thus, we may assume all $t_{B}$ to be distinct.
@@
  • The block of moves $m^{(i)}$ is legal and transfers $\ell_{B_{i}}+\ell^{\prime}_{B_{i}}$ coins from the pile to $B_{i}$, so
    $p_{i}\geq\ell_{B_{i}}+\ell^{\prime}_{B_{i}}$
@@
    $p_{i}\geq 2\ell_{B_{i}}-1$
  • $p_{i+1}=p_{i}-\ell_{B_{i}}$, so
    $2p_{i+1}$ $=2p_{i}-2\ell_{B_{i}}$
    $\geq 2p_{i}-(p_{i}+1)$
    $=p_{i}+1$
[…]
@@ 3.1 Definitions and main statement
We introduce the following definitions:
  • A (rectilinear) rectangle of dimensions $w\times h$ is a translate of $[0,w]\times[0,h]$.
  • A (rectilinear) diamond of dimensions $w\times h$ is a translate of the closed convex hull of the points (±$(\pm w/2,0),(0,\pm h/2)$. The points in a diamond with maximal and minimal $y$-coordinate are called the north/south poles, and the horizontal line segment passing through its center of mas is called its equator.
  • Given a bounded measurable set $E\subseteq\mathbb{R}^{2}$ of positive Lebesgue measure and $c>0$, denote by $cE$ the convex set that is a translate of $\{cx\mid x\in E\}$ and has the same center of mass as $E$. .
  • Give a collection $\mathcal{E}$ of bounded convex sets and $c>0$, we say that $\mathcal{E}$ is $c$-sparse if for every $E,E^{\prime}\in\mathcal{E}$ with $E\neq E^{\prime}$, we have $E\cap cE^{\prime}=cE\cap E^{\prime}=\emptyset$
  • A polygonal graph $\gamma\subseteq\mathbb{R}$ is the graph of a piecewise linear function $f:\mathbb{R}\rightarrow\mathbb{R}$. If the absolute values of slopes of all line segments in $\gamma$ is $\leq\alpha$ we say that $\gamma$ has slope $\leq\alpha$ (equivalently, $f$ is $\alpha$-Lipschitz).
  • A polygonal path will mean a polygonal path in some (not necessarily standard) coordinate system
@@
Let $w_{1},\ldots,w_{N}>0$ and $h_{1},\ldots,h_{N}\geq 0$. If
$h_{n}$ $\gg w_{n-1},h_{n-1}\hfill\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for }n=1,\ldots,N-1$
$w_{n}$ $\gg h_{n},w_{n-1}\hfill\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ for }n=1,\ldots,N$
then the following holds: To every sequence $\mathcal{R}=(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})$, where $\mathcal{R}_{n}$ is a $80$-sparse collection of rectilinear rectangles of dimensions $w_{n}\times h_{n}$, we can associate a set $\operatorname{safe}(\mathcal{R})\subseteq\mathbb{R}^{2}$, the set of safe points relative to $\mathcal{R}$, such that
@@
When applying the theorem, we often identify $(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})$ with $\mathcal{R}=\bigcup_{n=1}^{N}\mathcal{R}_{n}$. There is no ambiguity in this, provided that all rectangles in $\mathcal{R}_{n}$ have dimensions $w_{n}\times h_{n}$, and that these dimensions are distinct for different $n$, since then $\mathcal{R}_{1},\ldots,\mathcal{R}_{N}$ can be reconstructed from $\bigcup_{n=1}^{N}\mathcal{R}_{n}$.
We shall apply the proposition later to collections of rectangles in arbitrary orthogonal coordinate system. Since this section is devoted to the proof of the proposition above, we work only with rectilinear rectangles (and diamonds), and therefore omit the qualifier “rectilinear” for convenience.
@@ 3.2 Jigsaws
A jigsaw is a non-negative piecewise linear function function supported on a bounded interval. An $\alpha$-jigsaw is a jigsaw of slope $\leq\alpha$ (equivalently, $\alpha$-Lipschitz).
@@
This is the $\alpha$-jigsaw supported on $[x_{0}-y_{0}/\alpha,x_{0}+y_{0}/\alpha]$. Its graph on this interval is an isosceles triangle whose base is the interval above, whose sides of slope ±$\pm\alpha$, and whose its peak at $p$.
@@
  • $f\leq g\leq f+\Delta_{(x_{0},y_{0}-f(x_{0})),\beta-\alpha}$
  • $g|_{\mathbb{R}\setminus I}=f|_{\mathbb{R}\setminus I}$
  • The slope of $g$ on $I$ is bounded by $\beta$
@@ 3.3 Double jigsaws
[…]
@@
By joining regions of the types $A(g),A(h)$ above one can form more complicated shapes (e.g. dimaonds):
@@
A double jigsaw $G\subseteq\mathbb{R}^{2}$ with equator $J$ is a set of the form $G=(A(g^{+})\cup A(-g^{-}))+p$ and an interval $J=I+p$, where
  • $p\in\mathbb{R}^{2}$
@@
  • $I$ is the minimal closed interval supporting $g^{+},g^{-}$.
@@ 3.4 The hull of collections of diamonds
Let $I=[-r,r]$ be a symmetric interval and $g^{+},g^{-}$ jigsaws supported on $I$, and let $G=A(g^{+})\cup A(-g^{-})$ denote the corresponding double jigsaw.
@@
    The operation $\ltimes$ is asymmetric, because the right-hand term must be a diamond, while the left can be a jigsaw. But for pairs of diamonds it is commutative, because the maximum operation is. This allows the operation to be extanded to pairs of jigsaws generated by diamonds, and it is comutative on this class.
@@
    Otherwise, choose a set $G\in\mathcal{H}_{n}$ with $d(D,G)\leq h_{n}$, delete $G$ from $\mathcal{H}$ and replace it with $G\ltimes D$.
@@
    If $D\in\mathcal{D}_{n}$ was added in step (a) of round $n$ of Definition 3.8, then in the final hull, the double jigsaw $J$ containing $D$ is properly contained in a rectangle of dimensions $(w_{n}+10w_{\leq n-1})\times(h_{n}+10h_{n-1})$, and hence in a rectangle of dimensions $2w_{n}\times 2h_{n}$
@@
    Then $p\in\mathbb{R}^{2}\setminus\mathcal{H}^{\prime}$, where $\mathcal{H}^{\prime}$ is the hull of $\mathcal{D}^{\prime}_{1},\ldots,\mathcal{D}^{\prime}_{N}$
@@
(1) First, By induction, when the construction enters stage $n$, every $J\in\mathcal{H}$ has slopes at most $\alpha_{n+1}$ (or 00 when $n=N$).
For each $n$, we induct on the sequence $(D_{i})$ of elements of $\mathcal{D}_{n}$ processed in stage $n$ of the construction. Consider what happens when we reach $D_{i}$, and its distance from some double jigsaw $J\in\mathcal{H}$ is less than $h_{n}$. Let $(x_{0},y_{0})$ denote the center of $D_{i}$. Observe that $J$ arose by adjoining zero or more of the diamonds $D_{1},\ldots,D_{i-1}$ to a double jigsaw $J^{\prime}$ that was in $\mathcal{H}$ at the beginning of stage $n$, and, by our induction hypothesis, each time one of these $D_{j}$ was adjoined, the change to $J^{\prime}$ was limited to a box of dimensions $5w_{n}\times 5h_{n}$; since $\mathcal{D}_{n}$ is $20$-sparse, these changes are disjoint from the box of the same dimensions with center $(x_{0},y_{0})$. Thus, the jigsaws defining $J$ agree with those defining $J^{\prime}$ on the interval $[x_{0}-5w_{n},x_{0}+5w_{n}]$, and in particular its slopes on this interval are at most $\alpha_{n+1}$. This implies, in particular, that the closest point in $D_{i}$ to $J$ is its top or bottom vertex (or that one of these is actually contained in $J$), so the distance of the other vertex from the jigsaw must be at most $2h_{n}$. Now we invoke Lemma 3.3 to see that adjoining $D_{i}$ affects $J$ only in the interval $[x_{0}-r,x_{0}+r]$ for $r=2h_{n}/(\alpha_{n}-\alpha_{n+1})$, and since all changes happen in a region whose boundary has slopes $\alpha_{n}$, we have confined the change to a box of dimension $2r\times(\alpha_{n}\cdot 2r)$. Using the definition of $r$, and the assumptions $h_{n}=\alpha_{n}w_{n}$ and $\alpha_{n+1}<\alpha_{n}/10$, we see that the changes are confined to a box of dimensions $5w_{n}\times 5h_{n}$ containing $D_{i}$.
The second part of (1) follows from the first part and the sparsity assumption, because together they imply that the region added to the jigsaw when $D_{i}\in\mathcal{D}_{n}$ is added does not come within $h_{n}$ of any other diamond in $\mathcal{D}_{n}$, and so does not affect the choice of option (a) or (b) in future stages of the ocnstruction.
(2) follows because (1) shows that the additions in stage $n$ do not interact, and so if $J$ is a double jigsaw in $\mathcal{H}$ before entering stage $n$, then it is extended by at most $5h_{n}$ upwards and downwards, and $5w_{n}$ left and right, in the course of stage $n$. Each double jigsaw in the hull started from a diamond $D\in\mathcal{D}_{n}$ for some $n$, whose dimensions are $w_{n}\times h_{n}$. Thus the cumulative vertical height of a double jigsaw is $<h_{n}+10h_{\leq n-1}<2h_{n}$ and the cumulative width is $<w_{n}+10w_{\leq n-1}<2w_{n}$.

### v2

@@ 3 Coins and buckets (a combinatorial lemma)
Note that when coins are moved from one bucket to another we often do not control the orientation they will receive. In particular, after a coin is added to a non-empty bucket, that bucket is not oriented. Observe also that that legal moves do not change the total number of coins in a configuration, and that in each bucket the coins of a given orientation are interchangeable, in the sense that if one of them is to be moved, it does not matter which one it is.
@@
Let $c_{k,n}$ denote the configuration consisting of $k$ buckets, all of which are empty except for the first, which contains $n_{1}=n$ coins, out of which $n_{2}=\left\lfloor n_{1}/2\right\rfloor$ are heads and $\left\lceil n_{1}/2\right\rceil$ are tails. Consider the following sequence of moves starting from $c_{k,n}$. First, move $n_{2}$ heads from the first bucket to the second bucket, one by one, orienting them as “tails” whenever there is a choice. Thus, when the first coin lands in an empty bucket its orientation is tails, the second coin is forced to be heads, the next is again tails, and so on. When all $n_{2}$ coins have been transferred, we will have $n_{3}=\left\lfloor n_{2}/2\right\rfloor$ heads and $\left\lceil n_{2}/2\right\rceil$ tails in the second bucket. Now move all heads from the second bucket to the third in the same manner, and continue, ending in the $k$-th bucket. At the end $\left\lceil n_{i}/2\right\rceil$ tails remain in the buckets $i=1,\ldots,k-1$, and $n_{k}$ coins remain in the last bucket. The final configuration is oriented if and only if $n_{k}\leq 1$, and, tallying how many coins are left in each bucket, we have shown that $c_{k,n}$ is orientable if $n\leq 2^{k+1}-1$. Conversely,
@@
We begin the proof. Assume that $k,n$ are given and that $m_{1}\ldots m_{N}$ is a sequence of legal moves taking $c=c_{k,n}$ to an oriented configuration $c^{\prime}$. Our goal is to deduce that $n\leq 2^{k}-1$. Our strategy will be to reduce the given moves to a different sequence, in which each bucket is dealt with separately; then we will be able to directly analyze the situation similarly to the example above.
@@
To simplify analysis, we first modify the rules of the game. Introduce a special bucket called “the pile”, which serves as a temporary holding place for coins. While in the pile, coins do not have a specific orientation. We only allow moves that transfer a coin from a bucket to the pile, or from the pile to a bucket. The game begins with all coins in the pile, and must end in an empty pile and oriented buckets.
@@
    An original move that transfers a coin from bucket $B$ to bucket $B^{\prime}$ is replaced by two moves: one taking a coin from $B$ to the pile, and the other taking a coin from the pile to $B^{\prime}$.
Evidently, starting from $n$ coins on the pile and $k$ empty buckets, the first $n$ moves in the updated sequence result in the configuration $c$, and the remaining moves bring us to $c^{\prime}$.
@@
[…]
and let $\sigma_{B}$ denote their common orientation (if $B$ is empty in $c^{\prime}$ set $\sigma_{B}=$tails). Let $\sigma^{\prime}_{B}\neq\sigma_{B}$ denote the complementary orientation, and set
@@
Define the sequence of moves $\underline{m}_{B}=m_{B,1},\ldots,m_{B,\ell_{B}+2\ell^{\prime}_{B}}$ as follows:
  • The first $\ell_{B}+\ell^{\prime}_{B}$ moves insert coins into $B$, with the provision that if relevant, the orientation is $\sigma_{B}$.
@@
Then, if we apply $\underline{m}_{B}$ starting from a configuration in which $B$ is empty and there are at least $\ell_{B}+\ell^{\prime}_{B}$ coins in the pile, then the net result is that $\ell_{B}$ coins are transferred to $B$ with orientation $\sigma_{B}$. Note that no shorter sequence of moves can achieve this result.
@@
Let $t_{B}$ denote the maximal integer $1\leq t_{B}\leq N$ such that, after performing the move $m_{t_{B}}$, the bucket is in its peak state, but was not before. Observe that the move at time $t_{B}$, if it exists, must move a coin to $B$, so the times $t_{B}$ are distinct for different buckets, provided that $p_{B}>0$, while if $p_{B}=0$, then the bucket remains empty throughout, and we can delete all such buckets and reduce $k$. Thus, we may assume all $t_{B}$ to be distinct.
@@
  • The block of moves $m^{(i)}$ starts with a pile of size $p_{i-1}$, and it is legal and that transfers $\ell_{B_{i}}+\ell^{\prime}_{B_{i}}$ coins from the pile to $B_{i}$, so
    $p_{i-1}\geq\ell_{B_{i}}+\ell^{\prime}_{B_{i}}$
@@
    $p_{i-1}\geq 2\ell_{B_{i}}-1$
  • $p_{i}=p_{i-1}-\ell_{B_{i}}$, so
    $2p_{i}$ $=2p_{i-1}-2\ell_{B_{i}}$
    $\geq 2p_{i-1}-(p_{i-1}+1)$
    $=p_{i-1}-1$
  • $\ell_{k}\leq 1$ and $p_{k}=0$, so $p_{k-1}\leq 1$. Using the relation $p_{i}\leq 2p_{i+1}+1$ from the previous step, it follows inductively that
@@ 4.1 Definitions and main statement
We introduce the following definitions. All shapes are in $\mathbb{R}^{2}$ and widths, heights etc., are real numbers.
  • A (rectilinear) rectangle of dimensions $w\times h$ is a translate of $0,w]\times[0,h]$.
  • A (rectilinear) diamond of dimensions $w\times h$ is a translate of the closed convex hull of the points $(\pm w/2,0),(0,\pm h/2)$. The points in a diamond with maximal and minimal $y$-coordinate are called the north and south poles, and the horizontal line segment passing through its center of mass is called its equator.
  • Given a bounded measurable set $E\subseteq\mathbb{R}^{2}$ of positive Lebesgue measure and $c>0$, denote by $cE$ the convex set that is a translate of $\{cx\mid x\in E\}$ and has the same center of mass as $E$.
  • Given a collection $\mathcal{E}$ of bounded convex sets and $c>0$, we say that $\mathcal{E}$ is $c$-sparse if for every $E,E^{\prime}\in\mathcal{E}$ with $E\neq E^{\prime}$, we have $E\cap cE^{\prime}=cE\cap E^{\prime}=\emptyset$ .
  • A polygonal graph $\gamma\subseteq\mathbb{R}^{2}$ is the graph of a piecewise linear function $f:\mathbb{R}\rightarrow\mathbb{R}$. If the absolute values of slopes of all line segments in $\gamma$ is $\leq\alpha$ we say that $\gamma$ has slope $\leq\alpha$ (equivalently, $f$ is $\alpha$-Lipschitz).
  • A polygonal path will mean a polygonal graph in some (not necessarily standard) coordinate system
@@
Let $w_{1},\ldots,w_{N}>0$ and $h_{1},\ldots,h_{N}\geq 0$. Also denote $h_{0}=w_{0}=0$. If
$h_{n}$ $\gg w_{n-1},h_{n-1}\hfill\text{~~~ for }n=1,\ldots,N$
$w_{n}$ $\gg h_{n},w_{n-1}\hfill\text{~~~ for }n=1,\ldots,N$
then the following holds: To every sequence $\mathcal{R}=(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})$, where each $\mathcal{R}_{n}$ is an $80$-sparse collection of rectilinear rectangles of dimensions $w_{n}\times h_{n}$, we can associate a set $\operatorname{safe}(\mathcal{R})\subseteq\mathbb{R}^{2}$, the set of safe points relative to $\mathcal{R}$, such that
@@
When applying the theorem, we often identify $(\mathcal{R}_{1},\ldots,\mathcal{R}_{N})$ with $\mathcal{R}=\bigcup_{n=1}^{N}\mathcal{R}_{n}$. This is an abuse of notation since the latter does not always determine the former. If one think of the union as a multiset (with points having multiplicity counting how many times they appear in the union) then, with $w_{h},h_{n}$ given, there is no ambiguity.
We shall apply the proposition later to collections of rectangles in arbitrary orthonormal coordinate systems. Since this section is devoted to the proof of the proposition above, we work only with rectilinear rectangles (and diamonds), and therefore omit the qualifier “rectilinear” for convenience.
@@ 4.2 Jigsaws
A jigsaw is a non-negative piecewise linear function whose domain is a bounded interval in $\mathbb{R}$. An $\alpha$-jigsaw is a jigsaw of slope $\leq\alpha$ (equivalently, $\alpha$-Lipschitz).
@@
This is the $\alpha$-jigsaw supported on $x_{0}-y_{0}/\alpha,x_{0}+y_{0}/\alpha]$. Its graph on this interval is an isosceles triangle whose base is the interval above, whose sides are of slope $\pm\alpha$, and whose peak is at $p$.
@@
  • $f\leq g\leq f+\Delta_{(x_{0},y_{0}-f(x_{0})),\beta-\alpha}$ .
  • $g|_{\mathbb{R}\setminus I}=f|_{\mathbb{R}\setminus I}$ .
  • The slope of $g$ on $I$ is bounded by $\beta$ .
@@ 4.3 Double jigsaws
$A(g)=\{(x,y)\mid x\in I\;,\;0\leq y\leq g(x)\}$
If $h$ is a function such that h-h is a jigsaw, define $A(h)$ similarly:
@@
By joining regions of the types $A(g),A(h)$ above one can form more complicated shapes (e.g. diamonds):
@@
A double jigsaw $G\subseteq\mathbb{R}^{2}$ with equator $J\subseteq\mathbb{R}^{2}$ is a set of the form $G=(A(g^{+})\cup A(-g^{-}))+p$ and an interval $J=I+p$, where
  • $p\in\mathbb{R}^{2}$ .
@@
  • II is the minimal closed interval supporting $g^{+},g^{-}$.
@@ 4.4 The hull of collections of diamonds
Let I=[r,r]I=[-r,r] be a symmetric interval and $g^{+},g^{-}$ jigsaws supported on $I$, and let $G=A(g^{+})\cup A(-g^{-})$ denote the corresponding double jigsaw.
@@
    The operation $\ltimes$ is asymmetric, because the right-hand term must be a diamond, while the left can be a jigsaw. Even when both arguments are diamonds, it is not commutative.
@@
    Otherwise, choose a set $G\in\mathcal{H}$ with $d(D,G)\leq h_{n}$, delete $G$ from $\mathcal{H}$ and replace it with $G\ltimes D$.
@@
    If $D\in\mathcal{D}_{n}$ was added in step (a) of round $n$ of Definition 4.8, then in the final hull, the double jigsaw $J$ containing $D$ is properly contained in a rectangle of dimensions $(w_{n}+10w_{\leq n-1})\times(h_{n}+10h_{n-1})$, and hence in a rectangle of dimensions $2w_{n}\times 2h_{n}$ .
@@
    Then $p\in\mathbb{R}^{2}\setminus\mathcal{H}^{\prime}$, where $\mathcal{H}^{\prime}$ is the hull of $\mathcal{D}^{\prime}_{1},\ldots,\mathcal{D}^{\prime}_{N}$ .
@@
(1) First, By induction, when the construction enters stage $n$, every $J\in\mathcal{H}$ has slopes at most $\alpha_{n+1}$ (or 0 when $n=N$).
For each $n$, we induct on the sequence $(D_{i})$ of elements of $\mathcal{D}_{n}$ processed in stage $n$ of the construction. Consider what happens when we reach $D_{i}$, and its distance from some double jigsaw $J\in\mathcal{H}$ is less than $h_{n}$. Let $(x_{0},y_{0})$ denote the center of $D_{i}$. Observe that $J$ arose by adjoining zero or more of the diamonds $D_{1},\ldots,D_{i-1}$ to a double jigsaw $J^{\prime}$ that was in $\mathcal{H}$ at the beginning of stage $n$, and, by our induction hypothesis, each time one of these $D_{j}$ was adjoined, the change to $J^{\prime}$ was limited to a box of dimensions $5w_{n}\times 5h_{n}$; since $\mathcal{D}_{n}$ is $20$-sparse, these changes are disjoint from the box of the same dimensions with center $(x_{0},y_{0})$. Thus, the jigsaws defining $J$ agree with those defining $J^{\prime}$ on the interval $x_{0}-5w_{n},x_{0}+5w_{n}]$, and in particular its slopes on this interval are at most $\alpha_{n+1}$. This implies, in particular, that the closest point in $D_{i}$ to $J$ is its top or bottom vertex, or that these points are on different sides of the (line extending the) equator of $J$. Therefore, the distance of the other vertex from the jigsaw, or the equatorial line, is at most $2h_{n}$. Now we invoke Lemma 4.3 to see that adjoining $D_{i}$ affects $J$ only in the interval $x_{0}-r,x_{0}+r]$ for $r=2h_{n}/(\alpha_{n}-\alpha_{n+1})$, and since all changes happen in a region whose boundary has slopes $\alpha_{n}$, we have confined the change to a box of dimension $2r\times(\alpha_{n}\cdot 2r)$. Using the definition of $r$, and the assumptions $h_{n}=\alpha_{n}w_{n}$ and $\alpha_{n+1}<\alpha_{n}/10$, we see that the changes are confined to a box of dimensions $5w_{n}\times 5h_{n}$ containing $D_{i}$.
The second part of (1) follows from the first part and the sparsity assumption, because together they imply that the region added to the jigsaw when $D_{i}\in\mathcal{D}_{n}$ is added does not come within $h_{n}$ of any other diamond in $\mathcal{D}_{n}$, and so does not affect the choice of option (a) or (b) in future stages of the construction.
(2) follows because (1) shows that the additions in stage $n$ do not interact, and so if $J$ is a double jigsaw in $\mathcal{H}$ before entering stage $n$, then it is extended by at most $5h_{n}$ upwards and downwards, and $5w_{n}$ left and right, in the course of stage $n$. Each double jigsaw in the hull started from a diamond $D\in\mathcal{D}_{n}$ for some $n$, whose dimensions are $w_{n}\times h_{n}$. Thus the cumulative vertical height of a double jigsaw is $h_{n}+10h_{\leq n-1}<2h_{n}$ and the cumulative width is $w_{n}+10w_{\leq n-1}<2w_{n}$.